If the​ pre-tax cost function for​ John's Shoe Repair is:

C(q)=200+20q-2q2+0.333q3

and it faces a specific tax of

τ=​$20​,

what is the​ profit-maximizing condition if the market price is​ p? Can you solve for a​ single, profit-maximizing q in terms of​ p?

The​ profit-maximizing quantity in terms of p

q=
Answer #1

a at Eqm π Maxn P=MC Mc= 20-49 +.333 (3) q² +20. 80 P = 40-49 + .999q2 And Now P-40= .9999² -4q. R40 = 1.9999 - 28.99919.2_+(